OpenStudy (anonymous):

Write the equation of a line with the given information in point-slope form, through points (1,2) and (3,8).

OpenStudy (saifoo.khan):

find the slope first. 8-2 --- 3-1 6/2 3

OpenStudy (anonymous):

So here you have the same type problem as the last one, except that they didn't give you the slope. Big meenies! Fortunately you can find it using the slope formula.

OpenStudy (anonymous):

As stated above, the slope works out to be 3. So now plug that slope, along with either one of your two points into the point-slope formula for the line to get the equation.

OpenStudy (anonymous):

Slope is different than point-slope, correct?

OpenStudy (anonymous):

Yes. The slope formula uses two points to compute the slope between them: \[Slope = \frac{y_2 - y_1}{x_2-x_1}\]

OpenStudy (anonymous):

Saifoo, she's asked to put it into point-slope, no further work needed.

OpenStudy (saifoo.khan):

Yes, slope is one part of the equation. as in y = mx + c so "m" is the slope.

OpenStudy (saifoo.khan):

i think i did the same! :P

OpenStudy (anonymous):

Once you get the slope of the line (using the slope formula above) you just plug it into point-slope formula.

OpenStudy (anonymous):

You solved for slope-intercept form.

OpenStudy (anonymous):

Okay, i got the slope. what is slope-intercept?

OpenStudy (saifoo.khan):

BOTH ARE THE SAME THINGS MAN! >.<

OpenStudy (anonymous):

It's the : y = mx + b form of the equation.

OpenStudy (anonymous):

Stop shouting. If the question asks for point-slope, solving for slope-intercept is: a) more work b) likely to get her marked wrong

OpenStudy (anonymous):

You need to find the point-slope equation for the line, so plug it into the formula: y-y1 = m(x-x1)

OpenStudy (saifoo.khan):

OK, tell me the difference.. And i did from the same method!

OpenStudy (anonymous):

What set of ( ) do i put into the point-slope equation?

OpenStudy (anonymous):

Either point will work

OpenStudy (anonymous):

You can pick any point on that line and it'll work.

OpenStudy (saifoo.khan):

TELL ME THE DIFFERENCE.

OpenStudy (saifoo.khan):

??

OpenStudy (anonymous):

So it would be y-1=3(x-2)?

OpenStudy (anonymous):

You have your x1 and y1 mixed up there.

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

(x1,y1) = (1,2)

OpenStudy (saifoo.khan):

@polpak, Answer me too!

OpenStudy (anonymous):

Saifoo, stop spamming.

OpenStudy (saifoo.khan):

IM NOT SPAMMING! i did the same thing! why u deleted my answer again? i even showed the working.

OpenStudy (anonymous):

How do you delete someone elses answers?! Woaah...I didnt know you could do that.

OpenStudy (anonymous):

Did you get your answer victoria?

OpenStudy (saifoo.khan):

Please answer mee tooo!

OpenStudy (anonymous):

Yes, i did. Thankyou. (: Now only 8 more problems to go and im done woth all my summer homework! :D

OpenStudy (anonymous):

I am not going to argue with you Saifoo. I have neither the time nor the inclination.

OpenStudy (anonymous):

Good luck rather.

OpenStudy (saifoo.khan):

but wht was wrong with my answer answer?

OpenStudy (saifoo.khan):

lol, cancel one answer.

OpenStudy (saifoo.khan):

this is rude Polpak!

OpenStudy (anonymous):

I think your answer was fine. I didnt really understand it, but i dont seee what the problem is, you just did it a different way, and added one more step, by putting it into slope-intercept form. (:

OpenStudy (saifoo.khan):

He always does the same thing, deletes my answers and then answer himself. LOL.